Calculator

Pricing Strategy, Sales Variances

Eastman, Inc., manufactures and sells three products: R, S, and T. In January, Eastman, Inc., budgeted sales of the following.

Budgeted
Volume
Budgeted
Price
Product R 108,800        $25       
Product S 161,700        20       
Product T 23,700        20       

At the end of the year, actual sales revenue for Product R and Product S was $2,580,600 and $3,296,500, respectively. The actual price charged for Product R was $23 and for Product S was $19. Only $9 was charged for Product T to encourage more consumers to buy it, and actual sales revenue equaled $565,650 for this product.

Required:

1. Calculate the sales price and sales volume variances for each of the three products based on the original budget.

Sales price variance Sales volume variance
Product R $ $
Product S $ $
Product T $ $

2. Suppose that Product T is a new product just introduced during the year. What pricing strategy is Eastman, Inc., following for this product?
Answer #1
Sales Price Variance Sales Volume Variance
Product R 224400 UF 85000F
Product S 173500UF 236000F
Product T 691350 UF 783000F

Detail working for your easy understanding

Sales Price Variance= (Actual price-Budgeted Price) X Actual Quantity Sold
Product R= (23-25)X 112200 =$224400 UF
Actual Quantity Sold= 2580600/23=112200 Unit
Product S= (19-20)X 173500 =$173500 UF
Actual Quantity Sold= 3296500/19=173500 Unit
Product T= (9-20)X 62850 =$691350 UF
Actual Quantity Sold= 565650/9=62850 Unit
Sales Volume Variance=( Actual Sales Quantity- Budgeted Sales Quantity)X Budgeted Price
Product R= ( 112200-108800)X25=$85000 F
Product S= ( 173500-161700)X20=$236000 F
Product R= ( 62850-23700)X20=$783000 F

Part-2

Penetration pricing Policy to be followed for Product T